OpenStudy (anonymous):

a crew rowed 20 miles downstream, with the current, in 2 hours. the return trip upstream, against the current, same distance, took 5 hours. find the crews rowing rate and rate of the current

OpenStudy (anonymous):

i know i do rate * time =distance

OpenStudy (lgbasallote):

ahh downstream-upstream... downstream speed = speed of boat (still) + speed of current upstream speed = speed of boat (still) - speed of current let c = current; b = speed of boat (still) we need to find both huh...okay... it says downstream..they went 20 miles in 2 hours so...that's 10 mile/hr upstream speed...20 miles in 5 hrs so 4 mile/hr are you following? :)

OpenStudy (anonymous):

so.... R*2 = 20 R=10..... and then R=4 for the other, but idk exactly what i just solved as far as the rates?

OpenStudy (anonymous):

ya i am so far i feel like this is like the one the other day, so easy its hard :(

OpenStudy (lgbasallote):

hahaha no....look at the equation i wrote above.. downstream speed = speed of boat + current upstream speed = speed of boat - current we already have downstream and upstream speeds..and if you look close if you do a system there..current will calncel out and leave you with the speed of boat only right?

OpenStudy (lgbasallote):

current will cancel out*

OpenStudy (lgbasallote):

so...do you have an idea of the next step now? :D

OpenStudy (anonymous):

6?

OpenStudy (lgbasallote):

6? do you mean that's your speed of boat? i dont think so haha...why dont you post your solution here so we can see where you went wrong...

OpenStudy (anonymous):

so if the currents cancel out wouldnt that be 10-6

OpenStudy (anonymous):

10 being rate w current, and 4 against?

OpenStudy (lgbasallote):

no no no....we do a systems..okay i will do the next step.... i said we let b = boat; c = current so... 10 = b + c <---downstream speed 4 = b - c <----upstream speed do you see it now?

OpenStudy (anonymous):

how do i figure out the current speed then?

OpenStudy (anonymous):

33 ?s and i have probelm with one haha!! wth!

OpenStudy (anonymous):

wait so b=4 and 10? do i add them or subtract? i would assume subtract....which would be 6!

OpenStudy (anonymous):

ok ok... boat is 7 and current is 3???!!!

OpenStudy (lgbasallote):

no no no... 4 is the upstream speed..10 is the downstream speed..you do not have a b..you're looking for it remember? wait...do you know how to use systems of linear equations?

OpenStudy (anonymous):

not exactly i guess?

OpenStudy (anonymous):

i feel like the boat is 7 and current is 3

OpenStudy (anonymous):

actually i only have 3 minutes to answer this one ? so im going to just have to accept a missed one lol BUUTTTT thanks for all your help with this one

OpenStudy (lgbasallote):

hmmm how did you do it?

OpenStudy (lgbasallote):

actually...you're right :P

OpenStudy (anonymous):

i used substitution method actually... is that right then

OpenStudy (anonymous):

ok then boat is the row speed correct?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

THANK U SOOOO MUCH!!!

OpenStudy (lgbasallote):

<tips hat>